OpenStudy (anonymous):

You plug a string of 100 lights in series into a 120 V power outlet, and each light has a resistance of 3.00 Ω. If each light has a power rating of 0.50 W, what will happen? A. All the lights will go out. B. Only one light will go out. C. The string will remain lit. D. You cannot predict what will happen.

OpenStudy (michele_laino):

the total resistance is: 3*100=...ohms

OpenStudy (anonymous):

300!

OpenStudy (michele_laino):

ok! so the current is: I=120/300=...amperes

OpenStudy (anonymous):

0.4!

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

now the maximum current which can flow into each of the bulbs, is: I=W/R^2= 0.5/9=...amperes

OpenStudy (anonymous):

0.0555555...

OpenStudy (anonymous):

what does that mean? the lights will remain lit? :/

OpenStudy (michele_laino):

so, since 0.4>0.055, namely the current provided by the battery is greater than the maximum current which can flow into each bulb, wha can you conclude?

OpenStudy (anonymous):

ermm the string of lights will remain lit?

OpenStudy (anonymous):

choiceA?

OpenStudy (michele_laino):

sorry, I have made an error, the maximum current which can flow into each of the bulbs, is: \[I = \sqrt {\frac{W}{R}} = \sqrt {\frac{{0.5}}{3}} = ...\]

OpenStudy (anonymous):

0.408?

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

so, since 0.4<0.408, namely the current provided by the battery is less than the current which can flow into each bulb, what can you conclude?

OpenStudy (anonymous):

that means that all the lights will go out? choice A?

OpenStudy (michele_laino):

yes! that's right!

OpenStudy (anonymous):

yay!! thanks!!:D

OpenStudy (michele_laino):

:)